OpenStudy (anonymous):

Solve the following exponential equation. 36= 3^x+5 + 3^x+4

OpenStudy (anonymous):

\[36= 3^{x+5} + 3^{x+4}\] Divide everything by 3: \[12= {x+5} + {x+4}\] Combine like terms: \[12= 2x+9\] Solve for x: \[3= 2x\] \[1.5=x\]

OpenStudy (anonymous):

I think x = -2

OpenStudy (anonymous):

HI :D

OpenStudy (anonymous):

Hi...here's why...itll take a minute

OpenStudy (anonymous):

Yeah, x does =-2 Thanks anyways @Urebaa

OpenStudy (anonymous):

\[36 = 3^{4+x} + 3^{ 5+x}\] \[y = 3^{x; }; 3^{x+4}+3^{x+5}\] \[324 *3^{x} = 324 y = 36\] \[324y = 36 y = 19 3^{x = 1/9 3^{2+x}}=1\] \[3^{2+x}= 3^{0} 2+x = 0 x=-2\]

OpenStudy (anonymous):

\[36= 3^{x}\times3^{5} + 3^{x}\times3^{4}\] \[36 = 3^{x} (3^{5} + 3^{4} + 1)\] \[36 = 3^{x} {325)\] That's how i"ve been doing it so far

OpenStudy (anonymous):

The way i input got a little messed up.. sorry

OpenStudy (anonymous):

LOL yeah care to explain ?

OpenStudy (anonymous):

I'm confused as to what you're doing in your solution...

OpenStudy (anonymous):

Let me retype again...this time correctly.

OpenStudy (anonymous):

Sounds good. thanks :)

OpenStudy (anonymous):

Starting on the 3rd line...the first 2 were OK \[324 * 3^{x}=324\] \[y=36\] \[324y = 36\] \[y = 1/9\] \[3^{x}= 1/9\] \[3^{2+x}= 1\] \[3^{2+x}= 3^{0}\] 2+x = 0 x = -2

OpenStudy (campbell_st):

do it this way split the 1st power of 3 \[36 = 3 \times 3^{x + 4} + 3^{x + 4}\] so then factoring you get \[36 = 3^{x + 4}( 3 + 1)\] or \[36 = 3^{x +4}(4)\] divide both sides by 4 \[9 = 3^{x + 4}\] now rewrite 9 interms of a power of 3 \[3^2 = 3^{x + 4}\] equate the powers and solve for x 2 = x + 4

OpenStudy (campbell_st):

hope it makes sense...

OpenStudy (campbell_st):

and remember some of the basics for indices \[3^1 \times 3^{x +4} = 3^{1 + x + 4} = 3^{x + 5}\] when multiplying, add the powers

OpenStudy (anonymous):

I am still confused LOL out of what I wrong above, what am I doing wrong?

OpenStudy (anonymous):

I don't understand how either of you are solving the problem.

OpenStudy (anonymous):

wrote *

OpenStudy (campbell_st):

well you correctly showed that the terms can be split to show that \[36 = 3^x \times 3^5 + 3^x \times 3^4 \] you error is in factoring the correct factorisation is \[36 = 3^x(3^5 + 3^4)\] there is now + 1 so then the next step is division \[\frac{36}{3^5 + 3^4} = 3^x\] or \[\frac{36}{243 + 81} = 3^x\] which becomes \[\frac{36}{324} = 3^x\] simplifying and you get \[\frac{1}{9} = 3^x\] now write your fraction as a power of 3 \[3^{-2} = 3^x\] hence your answer. taking a larger common factor \[3^{x + 4}....instead .. of ... 3^x\] would have reduced the amount of work needed for the solution. so the error was early in your solution, ad was caused by an incorrect factorisation or an expression

OpenStudy (campbell_st):

oops should be there is no + 1

OpenStudy (anonymous):

WAIT , i get it. Thank you @campbell_st , your method explains it :) thanks for taking the time to explain it :)

OpenStudy (campbell_st):

glad to help.... your method does work with a little more care with factoring

OpenStudy (anonymous):

https://www.mathway.com this site doesn't help explaining it but it does give the answers lol @lovethecreation

OpenStudy (anonymous):

Thanks @tswavy :)